# From loop quantum gravity to cosmology: the 2-vertex model

### 22 Mar 2024 v1

@@ Abstract
We show that the U$(N)$-symmetry reduction of the loop-quantum-gravity truncation known as the 2-vertex model provides the effective improved dynamics of loop quantum cosmology with an arbitrary perfect barotropic fluid content.
@@ I Introduction
There are no closed general formulas to compute the volume of these polyhedra (apart from the tetrahedron). Therefore, we need to resort to numerical methods [25, 26] or to approximations to the volume formula. One such approximation is given in terms of the quadrupole moment for discrete surfaces, introduced in [27] and studied numerically in [17], whose determinant qualitatively replicates the behavior of the volume of the polyhedron. In this paper we calculate analytically the equation of motion for the volume approximated by the quadrupole moment.
The $\text{U}(N)$-symmetric sector of the 2-vertex model contains just one degree of freedom corresponding to the twist angle and its canonical conjugate momentum, i.e., the total area of the polyhedra. Comparing the evolution of the quadrupole approximation to the volume and the total area, we find that they precisely follow the evolution one expects from a homogeneous and isotropic evolution. We also find that a generalization of the 2-vertex Hamiltonian constraint (that preserves the symmetries of the reduced model) effectively takes into account matter content.
The central result of this paper is that the 2-vertex model is indeed an LQG truncation that effectively describes LQC with arbitrary barotropic perfect fluids. As a byproduct, we provide a clear interpretation of the improved dynamics parameter in terms of the 2-vertex main coupling.
@@ II.1 Spinorial formalism for LQG
The Poisson bracket between the holonomies and the fluxes is $\{g_{i},\vec{X}_{j}^{s,t}\}=\pm ig_{i}\vec{\sigma}\delta_{ij}/2$. It is invariant under the transformation (4). Furthermore, the Barbero-Immirzi parameter $\beta$ appears in these variables as it should according to LQG. The action of $g_{i}$ associates the (normalized) spinors on one vertex with the dual (normalized) spinors on the other vertex:
@@ IV.1 U($N$) symmetry
The Hamiltonian is invariant under the action of $\mathcal{E}_{ij}$. Thus initial conditions satisfying the U($N$) symmetry (that is, pairs of initial spinors satisfying  (22)) will ensure U($N$) symmetry throughout the whole evolution. Expressing the action on the phase space surface given by $\mathcal{E}_{ij}=0$ and applying the relation (22), we find that only one degree of freedom remains. If we choose that degree of freedom to be the angle $\varphi$, the canonical conjugate variable will be the total area of the polyhedra [14] $A=\sum_{i=1}^{N}X_{i}=\sum_{i=1}^{N}Y_{i}$, i.e. $\{\varphi,A\}=1$.
@@ V LQC from the U$(N)$-invariant 2-vertex model
In this section we show that the U$(N)$-invariant 2-vertex model is indeed a reduction of the full LQG theory which provides LQC improved effective dynamics.
@@ V.1 The constraint equation
$\breve{C}=\frac{2\gamma}{\beta^{2}}a\bigg{(}\cos\frac{\beta\pi_{a}}{2a}-1\bigg{)}-\frac{\kappa}{\gamma}a+a^{3}f(a),$ (35)
with $\kappa=-2\gamma(\lambda+\gamma)/\beta^{2}$, we have conveniently redefined the Lagrange multiplier as $\mathcal{M}=\mathcal{N}/a^{3}$, and $f(a)=2g(a^{2}/\beta)/(\beta^{2}a^{2})$ has dimensions of inverse length. This Hamiltonian constraint (35) can be written in terms of $(a,\dot{a})$ by using the equation of motion for $a$,
@@ V.2 FLRW cosmology as the small-twist regime of the 2-vertex model
In order to compare both dynamical equations (40) and (42) on equal footing it is necessary to introduce a fiducial volume $\mathcal{V}_{\Sigma}$ with respect to the line element $\mathop{}\!\mathrm{d}\Sigma_{k}^{2}$ (see footnote 3). This factor naturally appears multiplying both $\rho$ and $1/G$ when considering the Hamiltonian for FLRW dynamics. In this comparison we then see that the small-twist regime of the U$(N)$-invariant 2-vertex model describes a FLRW cosmology. The cosmological parameters and energy density are determined by the parameters $\gamma$, $\kappa$, and the function $f$ that define the 2-vertex model: The gravitational constant is given by $G=3\gamma\mathcal{V}_{\Sigma}/(8\pi)$, the spatial curvature by $k=\kappa$, and the matter content is a barotropic fluid with density $\rho=f/\mathcal{V}_{\Sigma}$ and pressure determined by the continuity equation $p=-(a^{3}f)^{\prime}/(3\mathcal{V}_{\Sigma}a^{2})$, the prime denoting derivative with respect to $a$. As a simple example, the 2-vertex model with $f(a)=\sum_{n}f_{n}a^{-\alpha_{n}}$ describes a cosmology filled with a number of perfect fluids with equations of state $p_{n}=w_{n}\rho_{n}$, where $w_{n}=\alpha_{n}/3-1$.
As we saw in Sec. IV.3, $\varphi=\xi$ can be interpreted as extrinsic curvature given by $\xi_{i}^{2}=-\ell^{2}\det(\hat{\ell}_{i}^{a}K_{a}^{I}\sigma_{I})=\ell^{2}\hat{\ell}_{i}^{a}\hat{\ell}_{i}^{b}K_{ac}K^{c}_{b}$. The extrinsic curvature of the spatial slices of a FLRW universe is proportional to the spatial metric $q_{ab}$, i.e. $K_{ab}=q_{ab}\dot{a}/(a\mathcal{N})$. This implies that $\xi_{i}=\ell\dot{a}/(a\mathcal{N})$, i.e. that it is independent of the edge $i$ in agreement with the U$(N)$-invariance, which can be written as
@@ V.3 LQC effective dynamics
Now we see that equation (37) has just the form of the usual LQC effective FLRW dynamics. More concretely, in LQC, for a universe with a free scalar field and $k=0$, the semiclassical states (i.e. Gaussian coherent states sharply peaked at the classical FLRW trajectories in the low-curvature regime) follow the effective Friedmann’s equation [33]
$\frac{1}{\mathcal{N}^{2}}\frac{\dot{a}^{2}}{a^{2}}=\frac{8\pi G}{3}\rho\Big{(}1-\frac{\rho}{\rho_{\text{sup}}}\Big{)},$ (46)
@@
to lowest order in the Gaussian spread of the semiclassical state, where $\rho$ is the energy density of the scalar field, $\rho_{\text{sup}}=(8\pi G\beta^{2}\Delta/3)^{-1}$ is the maximum value of the energy density (reached at the bounce), and $\Delta$ is the area gap. Thus, equation (37) with $\kappa=0$ (as corresponds to vanishing spatial curvature) and $\gamma=8\pi G/(3\mathcal{V}_{\Sigma})$ as before, indeed provides the LQC effective dynamics. Furthermore it tells us that the area gap must satisfy
$\frac{\sqrt{\Delta}}{a}=\frac{1}{2\gamma}=\frac{\ell}{2\beta a}.$ (47)
From this relation we can extract two conclusions. The length scale squared $\ell$ determines the area gap $\Delta$ and these quantities being “physical” length and area are naturally related to the 2-vertex model coupling constant $\gamma$ by the metric factor $a$. This dependence on the metric is well known in LQC. The improved dynamics prescription defines the minimum fiducial length $\bar{\mu}=\sqrt{\Delta}/a$ for the holonomies [9]. Therefore, the U$(N)$-invariant 2-vertex model, which is described within the full LQG theory, provides a clear meaning to the otherwise unjustified (from the fundamental point of view) improved dynamics as the LQC preferred dynamical framework. Indeed, the minimum fiducial length prescribed by the LQC improved dynamics [9] is precisely,
$\bar{\mu}=\frac{1}{2\gamma}=\frac{\ell}{2\beta a}.$ (48)
We see that $\bar{\mu}$ is just proportional to the constant characteristic length $\gamma^{-1}$ of the 2-vertex model (and in this sense directly determined by LQG); but that written in terms of the length scale $\ell$, acquires a metric factor as expected.
@@ VI Conclusions
Finally, we have shown that a generalization of the 2-vertex Hamiltonian constraint that preserves the symmetries of the model effectively takes into account matter content. Most importantly—and this is the central result of this work—we have proved that the 2-vertex model is indeed an LQG truncation that effectively describes LQC with arbitrary barotropic perfect fluids. As a byproduct, we have been able to provide a clear interpretation of the improved dynamics parameter $\bar{\mu}$ in terms of the 2-vertex main coupling $\gamma$, which has an intrinsic LQG origin.
@@ Acknowledgements.
ÁC acknowledges financial support from IPARCOS Ayudas de Máster IPARCOS-UCM/2022 and from the Spanish Government fellowship FPU22/03222. This work was supported by the Basque Government Grant IT1628-22, and by the Grant PID2021-123226NB-I00 (funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”). This work was also supported by the Grant PID2020-118159GB-C44 (funded by MCIN/AEI/10.13039/501100011033).

### 05 Apr 2024 v2

@@ Abstract
We study the notion of volume and its dynamics in the loop-quantum-gravity truncation known as the 2-vertex model. We also show that its U($N$)-symmetry reduction provides the old effective dynamics of loop quantum cosmology with an arbitrary perfect barotropic fluid content. A suitable modification of the Poisson bracket structure of the U($N$)-symmetric model leads to the LQC improved dynamics.
@@ I Introduction
There are no closed general formulas to compute the volume of these polyhedra (apart from the tetrahedron). Therefore, we need to resort to numerical methods [25, 26] or to approximations to the volume formula. One such approximation is given in terms of the quadrupole moment for discrete surfaces, introduced in [27] and studied numerically in [17], whose determinant qualitatively replicates the behavior of the volume of the polyhedron. In this paper we analytically calculate the equation of motion for the volume approximated by the quadrupole moment.
The $\text{U}(N)$-symmetric sector of the 2-vertex model contains just one degree of freedom corresponding to the twist angle and its canonical conjugate momentum, i.e., the total area of the polyhedra. Comparing the evolution of the quadrupole approximation to the volume and the total area, we find that they precisely follow the evolution one expects from a homogeneous and isotropic evolution. We also find that a generalization of the 2-vertex Hamiltonian constraint in the form of the addition of a function of the area (that preserves the symmetries of the reduced model) effectively takes into account matter content and, furthermore, fully determines the equation of state.
Finally, we show that the 2-vertex model is indeed an LQG truncation that effectively describes the LQC old dynamics [28, 29] with arbitrary barotropic perfect fluids. Furthermore, a suitable modification of the Poisson bracket structure of the U($N$)-symmetric model leads to the LQC improved dynamics [9].
@@ II.1 Spinorial formalism for LQG
The Poisson bracket between the holonomies and the fluxes is
$\{g_{i},\vec{X}_{j}^{s,t}\}=\pm ig_{i}\vec{\sigma}\delta_{ij}/2.$ (7)
The Barbero-Immirzi parameter $\beta$ divides the flux and (non-trivially) modifies the phases of the holonomy, in such a way that the Poisson bracket is invariant under the transformation (4) as should in LQG. The action of $g_{i}$ associates the (normalized) spinors on one vertex with the dual (normalized) spinors on the other vertex:
@@ IV.1 U($N$) symmetry
The Hamiltonian is invariant under the action of $\mathcal{E}_{ij}$. Thus initial conditions satisfying the U($N$) symmetry (that is, pairs of initial spinors satisfying  (23)) will ensure U($N$) symmetry throughout the whole evolution. Expressing the action on the phase space surface given by $\mathcal{E}_{ij}=0$ and applying the relation (23), we find that only one degree of freedom remains. If we choose that degree of freedom to be the angle $\varphi$, the canonical conjugate variable will be the total area of the polyhedra [14] $A=\sum_{i=1}^{N}X_{i}=\sum_{i=1}^{N}Y_{i}$, i.e.
$\{\varphi,A\}=1.$ (24)
@@ V LQC from the U$(N)$-invariant 2-vertex model
In this section we show that the U$(N)$-invariant 2-vertex model is indeed a reduction of the full LQG theory which provides LQC effective dynamics.
@@ V.1 The constraint equation
$\breve{C}=-\frac{4\gamma}{\beta^{2}}a\sin^{2}\frac{\beta\pi_{a}}{4a}-\frac{\kappa}{\gamma}a+a^{3}f(a),$ (37)
with $\kappa=-2\gamma(\lambda+\gamma)/\beta^{2}$, we have conveniently redefined the Lagrange multiplier as $\mathcal{M}=\mathcal{N}/a^{3}$, and $f(a)=2(\beta a)^{-2}g(a^{2}/\beta)$ has dimensions of inverse length. This Hamiltonian constraint (37) can be written in terms of $(a,\dot{a})$ by using the equation of motion for $a$,
@@ V.2 FLRW cosmology as the small-twist regime of the 2-vertex model
In order to compare both dynamical equations (42) and (44) on equal footing it is necessary to introduce a fiducial volume $\mathcal{V}_{\Sigma}$ with respect to the line element $\mathop{}\!\mathrm{d}\Sigma_{k}^{2}$ (see footnote 3). This factor naturally appears multiplying both $\rho$ and $1/G$ when considering the Hamiltonian for FLRW dynamics. In this comparison we then see that the small-twist regime of the U$(N)$-invariant 2-vertex model describes a FLRW cosmology. The cosmological parameters and energy density are determined by the parameters $\gamma$, $\kappa$, and the function $f$ that define the 2-vertex model: The gravitational constant is given by $G=3\gamma\mathcal{V}_{\Sigma}/(8\pi)$, the spatial curvature by $k=\kappa$, and the matter content is a barotropic fluid with density $\rho=f/\mathcal{V}_{\Sigma}$ and pressure determined by the continuity equation $p=-(a^{3}f)^{\prime}/(3\mathcal{V}_{\Sigma}a^{2})$, the prime denoting derivative with respect to $a$.
As a simple example, the 2-vertex model with
$f(a)=\sum_{n}f_{n}a^{-\alpha_{n}}$ (46)
describes a cosmology filled with a number of perfect fluids with equations of state
$p_{n}=w_{n}\rho_{n},\qquad w_{n}=\alpha_{n}/3-1.$ (47)
As we saw in Sec. IV.3, $\varphi=\xi$ can be interpreted as extrinsic curvature given by
$\xi_{i}^{2}=-\ell^{2}\det(\hat{\ell}_{i}^{a}K_{a}^{I}\sigma_{I})=\ell^{2}\hat{\ell}_{i}^{a}\hat{\ell}_{i}^{b}K_{ac}K^{c}_{b}.$ (48)
The extrinsic curvature of the spatial slices of a FLRW universe is proportional to the spatial metric $q_{ab}$, i.e. $K_{ab}=q_{ab}\dot{a}/(\mathcal{N}a)$. This implies that $\xi_{i}=\ell\dot{a}/(a\mathcal{N})$, i.e. that it is independent of the edge $i$ in agreement with the U$(N)$-invariance, which can be written as
@@ V.3 LQC effective dynamics
In the old dynamics of LQC for a universe with a free scalar field and vanishing spatial curvature, the semiclassical states (i.e. Gaussian coherent states sharply peaked at the classical FLRW trajectories in the low-curvature regime) follow the effective Friedmann’s equation [35]
$\frac{1}{\mathcal{N}^{2}}\frac{\dot{a}^{2}}{a^{2}}=\frac{8\pi G}{3}\rho\Big{(}1-\frac{a^{2}\rho}{\rho_{\star}}\Big{)},$ (51)
@@
to lowest order in the Gaussian spread of the semiclassical state, where $\rho$ is the energy density of the scalar field, $\rho_{\star}=(8\pi G\beta^{2}\Delta/3)^{-1}$, and $\Delta$ is the (constant) area gap. The maximum value of the energy density (reached at the bounce) is $\rho_{\text{max}}=\sqrt{\rho_{\star}^{3}/\rho_{0}}$, where $\rho_{0}=\rho\big{|}_{a=1}$.
Comparing Eq. (39) with $\kappa=0$ (as corresponds to vanishing spatial curvature) and (51) we observe that the U($N$)-symmetric 2-vertex model describes the old effective dynamics of LQC with an area gap $\Delta=1/(2\gamma)^{2}$ and a minimum fiducial length $\mu_{0}=\sqrt{\Delta}=1/(2\gamma)$.
V.4 Towards LQC improved dynamics
The U($N$)-symmetric 2-vertex model describes the old effective dynamics of LQC. Nevertheless, we can try to implement the improved dynamics [36] in a similar way to that of LQC by suitably modifying the model.
To do so, let us go back to the relation (23) imposed by the U($N$) symmetry on the spinors. Note that, in the U($N$) sector, the Poisson structure (3) for the spinors determines the Poisson bracket between the twist angle $\varphi$ and the total area $A$, so that they are canonically conjugate variables. To introduce the improved dynamics, we need to change this structure so that $\varphi$ and $A$ are no longer canonically conjugate. One way of doing this is to change the Poisson bracket to $\{\varphi,\sqrt{A}\}=1$, while keeping the Hamiltonian (35). Another equivalent option, which is the one we will use in the following, is to introduce a new variable $\phi=\varphi/\sqrt{A}$ canonically conjugate to $A$, so that $\{\phi,A\}=1$. Then, the U($N$) relation (23) now becomes
$|w_{i}\rangle=e^{-i\phi\sqrt{A}}|\bar{z}_{i}\rangle.$ (52)
The spinor Poisson structure that leads to this modification is no longer (3), so this modified model is not a direct consequence of the 2-vertex model but of a suitably modified one, whose structure is beyond the scope of this paper. The Hamiltonian in terms of these new canonical variables $\phi$ and $A$ is
$\breve{H}=2\mathcal{M}A^{2}\gamma\big{[}-1+\cos(\sqrt{A}\phi)+g(A)\big{]},$ (53)
where we have set $\kappa=0$ as before for comparison with LQC.
If we now follow steps analogous to those of Sec. V.3, we obtain
$\frac{1}{N^{2}}\frac{\dot{a}^{2}}{a^{2}}=\gamma f(a)\Big{[}1-\frac{\beta^{3}}{4\gamma}f(a)\Big{]}.$ (54)
On the other hand, the improved dynamics formalism introduces in the holonomies a non-constant minimum fiducial length $\bar{\mu}=\sqrt{\Delta}/a$. The resulting effective LQC equation is [37]
$\frac{1}{\mathcal{N}^{2}}\frac{\dot{a}^{2}}{a^{2}}=\frac{8\pi G}{3}\rho\Big{(}1-\frac{\rho}{\rho_{\star}}\Big{)},$ (55)
where now the maximum energy density (reached at the bounce) is precisely $\rho_{\star}$ and, thus, independent of the initial energy density $\rho_{0}$. We see that the modified U($N$)-symmetric 2-vertex dynamics (54) provides the effective improved dynamics (55) with energy density $\rho=f/\mathcal{V}_{\Sigma}$, gravitational constant $G=3\gamma\mathcal{V}_{\Sigma}/(8\pi)$ and area gap $\Delta=\beta/(4\gamma^{2})$. Consequently, the $\bar{\mu}$ of the LQC improved dynamics can be written in terms of the modified 2-vertex parameters as $\bar{\mu}=\sqrt{\Delta}/a=\sqrt{\beta}/(2\gamma a)$.
As we have already mentioned, this result is a consequence of an ad hoc modification of the 2-vertex model (very much in the same way as the improved dynamics is introduced in LQC). Nevertheless, this modification is useful to further understand the underlying physics behind the imposition of the improved dynamics in LQC and its relation with full LQG.
@@ VI Conclusions
We have shown that a generalization of the 2-vertex Hamiltonian constraint (that preserves the symmetries of the model) consisting of the addition of a function of the area, effectively takes into account matter content in the form of arbitrary barotropic perfect fluids. In particular, we have shown that there is a direct relation between the equation of state and the specific function of the area introduced in the 2-vertex Hamiltonian.
Finally, we have proved that the 2-vertex model is indeed an LQG truncation that effectively describes LQC with arbitrary barotropic perfect fluids. The LQC flavor that derives directly from the 2-vertex model is the old dynamics. However, a suitable modification of the Poisson bracket structure so that the twist angle is canonically conjugate to the square root of the area (and not the area itself) leads to the LQC improved dynamics.
@@ Acknowledgements.
We would like to thank Etera R. Livine and Mercedes Martín-Benito for very enlightening conversations. ÁC acknowledges financial support from IPARCOS Ayudas de Máster IPARCOS-UCM/2022 and from the Spanish Government fellowship FPU22/03222. This work was supported by the Basque Government Grant IT1628-22, and by the Grant PID2021-123226NB-I00 (funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”). This work was also supported by the Grant PID2020-118159GB-C44 (funded by MCIN/AEI/10.13039/501100011033).